# Vitali Telezkia and Stefan Klumppa

### v1

@@
[Uncaptioned image] Patterns of active dipolar particles in external magnetic fields
Vitali Telezkia and Stefan Klumppa
[Uncaptioned image] Active particles with a (magnetic) dipole moment are of interest for steering self-propelled motion, but also result in novel collective effects due to their dipole-dipole interaction. Here systems of active dipolar particles are studied with Brownian dynamics simulations to systematically characterize the different patterns they form, specifically in the presence of an external (magnetic) field. The combination of three types of order - clustering, orientational alignment and chain formation - is used to classify the patterns observed in these systems. In the presence of an external field, oriented chains and bands are found to be dominant. These structures show some similarities with columnar cluster seen in (passive) ferrofluids and display columnar spacing and number of lanes per cluster that both decrease with increasing field strength.
footnotetext: a University of Göttingen, Institute for the Dynamics of Complex Systems, Friedrich-Hund-Platz 1, 37077 Göttingen, Germany. E-mail: stefan.klumpp@phys.uni-goettingen.de
1 Introduction
Systems of self-propelled particles as a prime example of active matter have been studied extensively for their complex collective behavior 1, 2, 3, both from the purely theoretical point of view where these systems provide examples for systems intrinsically out of equilibrium 4 and from an applied point of view where self-propelled particles are model systems for microrobots with biomedical and environmental applications 5. From the applications point of view, the remote control of active particles is an important feature. An important method for remote control is through magnetic fields, provided that the active particles are equipped with a magnetic dipole moment 5, 6. Active particles with a magnetic dipole moment can be found in nature, in the form of magnetotactic bacteria 6, 7, but can also be made fully synthetically, e.g. active magnetic colloids 8, 9 and various forms of magnetic propellers and microrobots 10, 11, and as biohybrid systems by magnetically functionalizing motile microorganisms 12, 13.
While a large body of work has explored the magnetic steering of active particles as well as magnetically driven propulsion mechanisms 6, the effects of the magnetic interactions, both among the active particles and with an external magnetic field, have received considerably less attention. A few studies have explored the standard two-dimensional active Brownian particle model with additional magnetic interactions based on magnetic moments in the same plane as the active motion. Liao et al. 14 have determined the states of active particles with dipolar interactions in the absence of an external field and shown various patterns, in particular aggregation into chains. Similarly, our own earlier work has addressed patterns in small systems with only a few particles in confinement and studied the emergence of chain and rings 15. These and additional structures were also found computationally as well as experimentally using centimeter-sized toy robots in a recent study with small particle numbers and harmonic confinement 16. Moreover, motility-induced phase separation was shown to be suppressed by the magnetic interactions 14, 17. Mixtures, in which dipolar particles are not active themselves, but interact with active particles, have also been studied and also show aggregation into chains, but also large-scale clustering of the particles 18, 19. Finally, magnetic interactions were shown to result in the clustering of magnetotactic bacteria in flow through narrow channels 20, 21.
In this study, we extend the previous work 14, 15 to study the patterns that emerge in systems of dipolar active particles in the presence of an external field. We use systematic computer simulations to study the ordered states in such system as a function of the density, field strength, dipole moment and activity. We distinguish up to eight different states based on three order parameters introduced by Liao et al. 14 that describe the competition between chaining, clustering and orientation. As a reference, we start with the case without the external field and study how a magnetic field modulates these states. We show that under strong fields, bands of oriented particles are formed that constitute active particle equivalents of columnar clusters know from passive ferrofluids 22, 23, 24.
2 Methods
2.1 Model
We consider spherical active Brownian particles in two dimensions that have a permanent dipole moment $\vec{\mu}_{i}=\mu\hat{\vec{e}}_{i}$ with the magnetic strength $\mu$ at the particle center. Here and in the following, the superscript ’hat’ denotes a unit vector. The magnetic moment is aligned with the orientation of the particle $\hat{\vec{e}}_{i}$, which defines the direction of self-propulsion with speed $v_{0}$. The equations of motion for a particle $i$ with position $\vec{r}_{i}$ and orientation $\hat{\vec{e}}_{i}$ are given by
$\dot{\vec{r}_{i}}$ $=v_{0}\hat{\vec{e}}_{i}+\frac{1}{\gamma_{\mathrm{T}}}\vec{F}_{i}+\sqrt{2D_{\mathrm{T}}}\vec{\xi}_{i}^{\mathrm{T}}$ (1)
$\dot{\hat{\vec{e}}}_{i}$ $=\frac{1}{\gamma_{\mathrm{R}}}\vec{\tau}_{i}\times\hat{\vec{e}}_{i}+\sqrt{2D_{\mathrm{R}}}\vec{\xi}_{i}^{\mathrm{R}}\times\hat{\vec{e}}_{i},$ (2)
where $\gamma_{\mathrm{T}}$ and $\gamma_{\mathrm{R}}$ are the translational and rotational drag coefficients, $\vec{\xi}_{i}^{\mathrm{T}}$ is the translational stochastic force, and $\vec{\xi}_{i}^{\mathrm{R}}$ is the rotational stochastic torque. Both are described by Gaussian white noise with zero mean
$\vec{\xi}^{\mathrm{T,R}}\right>$ $=0$
$\vec{\xi}^{\mathrm{T,R}}\!\left(t\right)\cdot\vec{\xi}^{\mathrm{T,R}}\!\left(t^{\prime}\right)\right>$ $=\vec{1}\delta\!\left(t-t^{\prime}\right).$
$D_{\mathrm{T}}$ and $D_{\mathrm{R}}$ are the corresponding diffusion coefficients. $\vec{F}_{i}$ and $\vec{\tau}_{i}$ describe forces and torques acting on particle $i$. The forces $\vec{F}_{i}$ acting on particle $i$ are given by the sum of the particle-particle interactions
$\vec{F}_{i}=\sum_{i\neq j}\vec{F}^{\mathrm{dd}}_{ij}+\sum_{i\neq j}\vec{F}^{\mathrm{WCA}}_{ij}$
The pairwise forces between two particles $i$, $j$ separated by the distance $r_{ij}=\left|\vec{r}_{ij}\right|=\left|\vec{r}_{i}-\vec{r}_{j}\right|$ consist of dipolar forces and excluded volume interactions. The dipolar forces are given by 25
$\vec{F}^{\mathrm{dd}}_{ij}=$ $\frac{3\mu_{0}\mu^{2}}{4\pi r_{ij}^{4}}\left[\hat{\vec{r}}_{ij}\left(\hat{\vec{e}}_{i}\cdot\hat{\vec{e}}_{j}\right)+\hat{\vec{e}}_{i}\left(\hat{\vec{r}}_{ij}\cdot\hat{\vec{e}}_{j}\right)+\hat{\vec{e}}_{j}\left(\hat{\vec{r}}_{ij}\cdot\hat{\vec{e}}_{i}\right)\right.$
[…]
@@ 2.1 Model
The excluded volume interactions are modeled by the Weeks-Chandler-Anderson (WCA) 26 pair potential with the interaction energy $\epsilon$ and the particle diameter $\sigma$
@@
$\vec{F}^{\mathrm{WCA}}_{ij}=\begin{cases}48\epsilon\frac{\vec{r}_{ij}}{r_{ij}^{2}}\left[\left(\frac{\sigma}{r_{ij}}\right)^{12}-\frac{1}{2}\left(\frac{\sigma}{r_{ij}}\right)^{6}\right]&\text{if $r_{ij}<2^{1/6}\,\sigma$}\\ 0&\text{if $r_{ij}\geq 2^{1/6}\,\sigma$}.\end{cases}$
@@
The torques $\vec{\tau}_{i}$ acting on the particle $i$ result from the pairwise dipolar interactions 27
$\vec{\tau}_{i}=\sum_{i\neq j}\vec{\tau}^{\mathrm{dd}}_{ij}=\sum_{i\neq j}\frac{\mu_{0}\mu^{2}}{4\pi r_{ij}^{3}}\left[3\left(\hat{\vec{e}}_{i}\cdot\hat{\vec{r}}_{ij}\right)\left(\hat{\vec{e}}_{j}\times\hat{\vec{r}}_{ij}\right)+\left(\hat{\vec{e}}_{i}\times\hat{\vec{e}}_{j}\right)\right].$
2.2 Simulation method
@@
The dynamics of active dipolar particles was simulated using the simulation toolkit HOOMD-blue, version2.6 28, 29. The active self-propulsion of the particles and their alignment in an external magnetic field were added through modules developed in-house. The project data and its parameter space was managed within a signac framework 30, 31.
@@
All simulations used $N=1156$ particles and periodic boundary conditions. To avoid interactions between periodic images of the particles, a cut-off distance $r_{\mathrm{cut}}=L/2-\sigma$ was introduced for dipolar interactions. To recover correct rotational diffusion dynamics for spheres, we set $\gamma_{R}=3\gamma_{T}$ within HOOMD-blue. To study the effect of density, we varied the side length of the simulation box. We chose the following three densities. One well below ($\Phi=0.13,\,L=85$), one close to ($\Phi=0.23,\,L=63$) and one above ($\Phi=0.57,\,L=40$) the critical density where MIPS is expected in systems of active Brownian particles without dipolar interactions. To investigate the effect of a constant homogeneous external field on the collective dynamics of dipolar active particles, we systematically explored different combinations of the strength $B$ of the external field and the magnetic moment of the particles.
@@
As in our previous work 15, we used the following dimensionless units: distances are measured in units of the particle size $\sigma$ and time in units of $\sigma^{2}/D_{0}^{T}$, the time in which a particle diffuses over a distance of its own size. The strength of the WCA interaction $\epsilon$ is used the unit of energy, thus temperature has units of $\epsilon/k_{B}$. All magnetic quantities are expressed relative to a reference field strength $B_{0}=1\times 10^{-5}$ T, which is of the order of the magnetic field of the Earth. Thus, the strength of the magnetic field has units of $B_{0}$ and magnetic moments are expressed units of $\epsilon/B_{0}$. Note that with these units, the self-propulsion velocity $v_{0}$ is identical to the Peclet number $v_{0}\sigma/D_{0}^{T}$.
2.3 Order parameters
@@
In large systems of active dipolar particles, complex structures such as branches, connected networks, bands or loops within rings and large scale cluster formation are expected to emerge 32, 32, 14, 18, 19, 33. To characterize such patterns, we used the following three global order parameters introduced by Liao et al. 14. These order parameters were averaged over the simulation trajectories from $t=18$ to $t=20$ with data taken every $\Delta t=0.1$. We used combinations of these three order parameters to identify and classify different state of the system and to construct state diagrams as described below.
@@
Similar to other active systems, we expect active dipolar particles to form large clusters. To quantify clustering formation, we determined the fraction of all particles that are part of the largest cluster by counting the number of particles $n^{*}_{\mathrm{c}}$ in that cluster relative to the total particle number of the system $N$,
@@
The fraction of the largest cluster is close to 0 when particles form small or no clusters and reaches 1 when the whole system forms one giant cluster. To identify clusters and, specifically, the largest cluster, we performed a density based cluster analysis (DBSCAN 34) with the value of the cut-off distance $\epsilon$ based on the nearest neighbor distance $g_{1}(r)$ for each corresponding density.
@@
$\phi_{\mathrm{e}}=\frac{1}{N}\left|\sum_{i}\hat{\vec{e}}_{i}\right|.$ (4)
Finally, to assess the degree of chain formation in large systems, we applied the following chain criteria, following Liao et al. 14. Two particles are considered to be bonded if they fulfill the following three criteria: First, their distance is smaller than the cut-off distance, $r_{ij}<\epsilon$. Second, they are aligned in parallel $\hat{\vec{e}}_{i}\cdot\hat{\vec{e}}_{j}>0$, and third their orientation is in line with their connecting distance vector $\left(\hat{\vec{e}}_{i}\times\hat{\vec{r}}_{ij}\right)\cdot\left(\hat{\vec{e}}_{j}\times\hat{\vec{r}}_{ij}\right)>0$. The fraction of particles that fulfill all three criteria give the degree of polymerization
@@ 3 Results
In this study, we investigated how a homogeneous external magnetic field affects structure formation and collective behavior of (magnetic) dipolar active particles. To this end, we performed extensive Brownian dynamics simulations of active dipolar particles in a homogeneous external magnetic field while systematically varying the field strength $B$, the active velocity $v_{0}$, and the magnetic moment $\mu$ (all in dimensionless units, see Methods). We performed our simulations for three representative densities $\Phi=$ 0.13, 0.23, 0.57. We classified the different states of the system with respect to their clustering, orientational order, and chain formation, summarizing the results in state diagrams.
3.1 No external field

### v2

# Patterns of active dipolar particles in external magnetic fields

@@
Abstract
Active particles with a (magnetic) dipole moment are of interest for steering self-propelled motion, but also result in novel collective effects due to their dipole-dipole interaction. Here systems of active dipolar particles are studied with Brownian dynamics simulations to systematically characterize the different patterns they form, specifically in the presence of an external (magnetic) field. The combination of three types of order - clustering, orientational alignment and chain formation - is used to classify the patterns observed in these systems. In the presence of an external field, oriented chains and bands are found to be dominant. These structures show some similarities with columnar cluster seen in (passive) ferrofluids and display columnar spacing and number of lanes per cluster that both decrease with increasing field strength.
I Introduction
Systems of self-propelled particles as a prime example of active matter have been studied extensively for their complex collective behavior [1, 2, 3], both from the purely theoretical point of view where these systems provide examples for systems intrinsically out of equilibrium [4] and from an applied point of view where self-propelled particles are model systems for microrobots with biomedical and environmental applications [5]. From the applications point of view, the remote control of active particles is an important feature. An important method for remote control is through magnetic fields, provided that the active particles are equipped with a magnetic dipole moment [5, 6]. Active particles with a magnetic dipole moment can be found in nature, in the form of magnetotactic bacteria [6, 7], but can also be made fully synthetically, e.g. active magnetic colloids [8, 9] and various forms of magnetic propellers and microrobots [10, 11], and as biohybrid systems by magnetically functionalizing motile microorganisms [12, 13].
While a large body of work has explored the magnetic steering of active particles as well as magnetically driven propulsion mechanisms [6], the effects of the magnetic interactions, both among the active particles and with an external magnetic field, have received considerably less attention. A few studies have explored the standard two-dimensional active Brownian particle model with additional magnetic interactions based on magnetic moments in the same plane as the active motion. Liao et al. [14] have determined the states of active particles with dipolar interactions in the absence of an external field and shown various patterns, in particular aggregation into chains. Similarly, our own earlier work has addressed patterns in small systems with only a few particles in confinement and studied the emergence of chain and rings [15]. These and additional structures were also found computationally as well as experimentally using centimeter-sized toy robots in a recent study with small particle numbers and harmonic confinement [16]. Moreover, motility-induced phase separation was shown to be suppressed by the magnetic interactions [14, 17]. Mixtures, in which dipolar particles are not active themselves, but interact with active particles, have also been studied and also show aggregation into chains, but also large-scale clustering of the particles [18, 19]. Finally, magnetic interactions were shown to result in the clustering of magnetotactic bacteria in flow through narrow channels [20, 21].
In this study, we extend the previous work [14, 15] to study the patterns that emerge in systems of dipolar active particles in the presence of an external field. We use systematic computer simulations to study the ordered states in such system as a function of the density, field strength, dipole moment and activity. We distinguish up to eight different states based on three order parameters introduced by Liao et al. [14] that describe the competition between chaining, clustering and orientation. As a reference, we start with the case without the external field and study how a magnetic field modulates these states. We show that under strong fields, bands of oriented particles are formed that constitute active particle equivalents of columnar clusters know from passive ferrofluids [22, 23, 24].
II Methods
II.1 Model
We consider spherical active Brownian particles in two dimensions that have a permanent dipole moment $\bm{\mu}_{i}=\mu\hat{\bm{e}}_{i}$ with the magnetic strength $\mu$ at the particle center. Here and in the following, the superscript ’hat’ denotes a unit vector. The magnetic moment is aligned with the orientation of the particle $\hat{\bm{e}}_{i}$, which defines the direction of self-propulsion with speed $v_{0}$ (fig. 1). The equations of motion for a particle $i$ with position $\bm{r}_{i}$ and orientation $\hat{\bm{e}}_{i}$ are given by
$\dot{\bm{r}_{i}}$ $=v_{0}\hat{\bm{e}}_{i}+\frac{1}{\gamma_{\mathrm{T}}}\bm{F}_{i}+\sqrt{2D_{\mathrm{T}}}\bm{\xi}_{i}^{\mathrm{T}}$ (1)
$\dot{\hat{\bm{e}}}_{i}$ $=\frac{1}{\gamma_{\mathrm{R}}}\bm{\tau}_{i}\times\hat{\bm{e}}_{i}+\sqrt{2D_{\mathrm{R}}}\bm{\xi}_{i}^{\mathrm{R}}\times\hat{\bm{e}}_{i},$ (2)
where $\gamma_{\mathrm{T}}$ and $\gamma_{\mathrm{R}}$ are the translational and rotational drag coefficients, $\bm{\xi}_{i}^{\mathrm{T}}$ is the translational stochastic force, and $\bm{\xi}_{i}^{\mathrm{R}}$ is the rotational stochastic torque. Both are described by Gaussian white noise with zero mean
$\xi^{\alpha}_{i}\!\left(t\right)\xi^{\beta}_{j}\!\left(t^{\prime}\right)\right>$ $=\delta_{i,j}\delta_{\alpha,\beta}\delta\!\left(t-t^{\prime}\right),$
where the upper index ($\alpha,\beta$) runs over the $x$- and $y$-components of the translational noise as well as the rotational noise ($z$ component), and the lower index ($i,j$) runs over the particles. $D_{\mathrm{T}}$ and $D_{\mathrm{R}}$ are the corresponding diffusion coefficients. $\bm{F}_{i}$ and $\bm{\tau}_{i}$ describe forces and torques acting on particle $i$. The forces $\bm{F}_{i}$ acting on particle $i$ are given by the sum of the particle-particle interactions
$\bm{F}_{i}=\sum_{i\neq j}\bm{F}^{\mathrm{dd}}_{ij}+\sum_{i\neq j}\bm{F}^{\mathrm{WCA}}_{ij}$
The pairwise forces between two particles $i$, $j$ separated by the distance $r_{ij}=\left|\bm{r}_{ij}\right|=\left|\bm{r}_{i}-\bm{r}_{j}\right|$ consist of dipolar forces and excluded volume interactions. The dipolar forces are given by [25]
$\bm{F}^{\mathrm{dd}}_{ij}=$ $\frac{3\mu_{0}\mu^{2}}{4\pi r_{ij}^{4}}\left[\hat{\bm{r}}_{ij}\left(\hat{\bm{e}}_{i}\cdot\hat{\bm{e}}_{j}\right)+\hat{\bm{e}}_{i}\left(\hat{\bm{r}}_{ij}\cdot\hat{\bm{e}}_{j}\right)+\hat{\bm{e}}_{j}\left(\hat{\bm{r}}_{ij}\cdot\hat{\bm{e}}_{i}\right)\right.$
@@ II.1 Model
The excluded volume interactions are modeled by the Weeks-Chandler-Anderson (WCA) [26] pair potential with the interaction energy $\epsilon$ and the particle diameter $\sigma$
@@
$\bm{F}^{\mathrm{WCA}}_{ij}=\begin{cases}48\epsilon\frac{\bm{r}_{ij}}{r_{ij}^{2}}\left[\left(\frac{\sigma}{r_{ij}}\right)^{12}-\frac{1}{2}\left(\frac{\sigma}{r_{ij}}\right)^{6}\right]&\text{if $r_{ij}<2^{1/6}\,\sigma$}\\ 0&\text{if $r_{ij}\geq 2^{1/6}\,\sigma$}.\end{cases}$
@@
The torques $\bm{\tau}_{i}$ acting on the particle $i$ result from the pairwise dipolar interactions [27] and the interaction with the external homogeneous magnetic field ${\bm{B}}=B\hat{\bm{e}}_{x}$ (stationary frame of reference),
$\bm{\tau}_{i}$ $=\sum_{i\neq j}\bm{\tau}^{\mathrm{dd}}_{ij}+\mu\hat{\bm{e}}_{i}\times{\bm{B}}$
$=\sum_{i\neq j}\frac{\mu_{0}\mu^{2}}{4\pi r_{ij}^{3}}\left[3\left(\hat{\bm{e}}_{i}\cdot\hat{\bm{r}}_{ij}\right)\left(\hat{\bm{e}}_{j}\times\hat{\bm{r}}_{ij}\right)+\left(\hat{\bm{e}}_{i}\times\hat{\bm{e}}_{j}\right)\right]+\mu\hat{\bm{e}}_{i}\times{\bm{B}}.$
Figure 1: Schematic illustrating single active particle with orientation $\hat{\bm{e}}$ in external homogeneous magnetic field $B\hat{\bm{e}}_{x}$ (blue). Particle has magnetic strength $\mu$ and self-propulsion speed $v_{0}$ (back).
II.2 Simulation method
@@
The dynamics of active dipolar particles were simulated using the simulation toolkit HOOMD-blue, version 2.6 [28, 29]. The active self-propulsion of the particles and their alignment in an external magnetic field were added through modules developed in-house. The project data and its parameter space was managed within a signac framework [30, 31].
@@
All simulations used $N=1156$ particles and periodic boundary conditions. This relatively small system size was chosen to keep the simulation time manageable while systematically scanning a rather large parameter space. The same system size was used in previous work without an external field, where it was also shown that cluster sizes saturated for $N>1000$ [14]. To avoid interactions between periodic images of the particles, a cut-off distance of the order of half the system size, $r_{\mathrm{cut}}=L/2-\sigma$, was introduced for dipolar interactions. To recover correct rotational diffusion dynamics for spheres, we set $\gamma_{R}=3\gamma_{T}$ within HOOMD-blue. To study the effect of density, we varied the side length of the simulation box. We chose the following three densities. One well below ($\Phi=$0.13$,\,L=$85$$), one close to ($\Phi=$0.23$,\,L=$63$$) and one above ($\Phi=$0.57$,\,L=$40$$) the critical density where MIPS is expected in systems of active Brownian particles without dipolar interactions. To investigate the effect of a constant homogeneous external field on the collective dynamics of dipolar active particles, we systematically explored different combinations of the strength $B$ of the external field and the magnetic moment of the particles.
@@
As in our previous work [15], we used the following dimensionless units: distances are measured in units of the particle size $\sigma$ and time in units of $\sigma^{2}/D_{0}^{T}$, the time in which a particle diffuses over a distance of its own size. The strength of the WCA interaction $\epsilon$ is used as the unit of energy, thus temperature $T$ has units of $\epsilon/k_{B}$ (here, $T=1$). All magnetic quantities are expressed relative to a reference field strength $B_{0}=1\times 10^{-5}$ T, which is of the order of the magnetic field of the Earth. Thus, the strength of the magnetic field has units of $B_{0}$ and magnetic moments are expressed units of $\epsilon/B_{0}$. Note that with these units, the self-propulsion velocity $v_{0}$ is identical to the Peclet number $v_{0}\sigma/D_{0}^{T}$. Numerical integration was performed for $N_{t}=10^{6}$ time steps with an integration time step width $\delta_{t}=2\times 10^{-5}$. Simulations were evaluated from $t=18$ to $t=20$ with data taken every $\Delta t=0.1$.
II.3 Order parameters
@@
In large systems of active dipolar particles, complex structures such as branches, connected networks, bands or loops within rings and large scale cluster formation are expected to emerge [32, 32, 14, 18, 19, 33]. To characterize such patterns, we used the following three global order parameters introduced by Liao et al. [14]. We used combinations of these three order parameters to identify and classify different state of the system and to construct state diagrams as described below.
@@
Similar to other active systems, we expect active dipolar particles to form large clusters. To quantify cluster formation, we determined the fraction of all particles that are part of the largest cluster by counting the number of particles $n^{*}_{\mathrm{c}}$ in that cluster relative to the total particle number of the system $N$,
@@
The fraction of the largest cluster is close to 0 when particles form small or no clusters and reaches 1 when the whole system forms one giant cluster. To identify clusters and, specifically, the largest cluster, we performed a density based cluster analysis (DBSCAN [34]) with the value of the cut-off distance $\delta_{c}$ based on the nearest neighbor distance $g_{1}(r)$ for each corresponding density.
@@
$\phi_{\mathrm{e}}=\frac{1}{N}\left|\sum_{i}\hat{\bm{e}}_{i}\right|.$ (4)
Finally, to assess the degree of chain formation in large systems, we applied the following chain criteria, following Liao et al. [14]. Two particles are considered to be bonded if they fulfill the following three criteria: First, their distance is smaller than the cut-off distance, $r_{ij}<\delta_{c}$. Second, they are aligned in parallel $\hat{\bm{e}}_{i}\cdot\hat{\bm{e}}_{j}>0$, and third their orientation is in line with their connecting distance vector $\left(\hat{\bm{e}}_{i}\times\hat{\bm{r}}_{ij}\right)\cdot\left(\hat{\bm{e}}_{j}\times\hat{\bm{r}}_{ij}\right)>0$. The fraction of particles that fulfill all three criteria give the degree of polymerization
@@ III Results
In this study, we investigated how a homogeneous external magnetic field affects structure formation and collective behavior of (magnetic) dipolar active particles. To this end, we performed extensive Brownian dynamics simulations of active dipolar particles in a homogeneous external magnetic field while systematically varying the field strength $B$, the active velocity $v_{0}$, and the magnetic moment $\mu$ (all in dimensionless units, see Methods). We performed our simulations for three representative densities $\Phi=$0.13 , 0.23 and 0.570.130.230.570.130.230.57start_ARG 0.13 end_ARG , start_ARG 0.23 end_ARG and start_ARG 0.57 end_ARG. We classified the different states of the system with respect to their clustering, orientational order, and chain formation, summarizing the results in state diagrams.
III.1 No external field